OpenStudy (anonymous):

Use the technique of completing the square to transform the quadratic equation into the form (x+c)2 = a. 2x2 – 28x + 84 = 0

OpenStudy (precal):

factor out the 2 from the first 2 terms

OpenStudy (zepp):

Let's first factor the 2 out of this: \(2(x^2=14x+42)=0\)

OpenStudy (zepp):

@precal you stole my line lol :P

OpenStudy (precal):

sorry, great minds think alike but you put an equal sign instead of a minus

OpenStudy (precal):

Listen to zepp

OpenStudy (anonymous):

How do you get rid of the two, though?

OpenStudy (zepp):

Then you need to take the middle term's coefficient, divide it by 2, square it and you'll get the constant that's supposed to be behind it so it becomes a perfect square.

OpenStudy (zepp):

You divide by 2 to both side and it would vanish! :D because 0/2=0

OpenStudy (zepp):

Oh lol @precal - and = are so close from each other :P

OpenStudy (precal):

yes but have 2 distinct meanings in our little world of math

OpenStudy (anonymous):

So I'm left with (x^2=14x+42) ?

OpenStudy (zepp):

\(2(x^2-14x+42)=0\\\frac{2(x^2-14x+42)}{2}=\frac{0}{2}\\(x^2-14x+42)=0\)

OpenStudy (zepp):

Exactly, now you are able to continue.

OpenStudy (anonymous):

Is there a 0 at the end?

OpenStudy (zepp):

Now take the coeffient of the middle term (with 1 power), which is 14,divide by 2, (7), and square it, 49, this would be the term you should have so you could complete the square :) Yes, it is an equation =)

OpenStudy (anonymous):

How did I end up with (x^2=49x+42)=0?

OpenStudy (zepp):

but we need 49, not 42! What are we going to do? Add 7 to get 49, but also don't forget to substract it \(x^2-14x+42+7-7=0\)

OpenStudy (zepp):

Or, \(x^2-14x+42+7=0+7\)

OpenStudy (anonymous):

x^2=14x+49=7?

OpenStudy (zepp):

Very good, although it's a -, not =

OpenStudy (anonymous):

-7 or x^2-?

OpenStudy (zepp):

Now you are all set, you may proceed :) x^2-

OpenStudy (anonymous):

How?

OpenStudy (zepp):

Well, you got a perfect square right the, the middle term's coefficient correspond to the half of the square root of the last term (constant)

OpenStudy (anonymous):

Do I end up with -12x+7?

OpenStudy (zepp):

\(x^2-14x+49=(x-7)^2\) :D

OpenStudy (zepp):

\(x^2-14x+49=7\\(x-7)^2=7\\(x-7)^2-7=0\)

OpenStudy (zepp):

Is that clear enough? If it's not, I can re-explain :)